# Learning to Use AI for Learning: How Can We Effectively Teach and Measure Prompting Literacy for K–12 Students?

### v1

@@ Introduction
Most learners (87%) reported that they learned AI-related knowledge, including how to use AI in learning (e.g. ”I can’t ask AI to write answers directly, otherwise I won’t learn anything.”), how to ask AI questions effectively (e.g. ”Adding background and context information in the prompt can help AI to form better answers”), and AI’s capabilities (e.g. ”AI can help on exam preparation”, ”AI also has problems that it doesnt́ know about”). These indicated learners’ gain on conceptual knowledge and skills of using AI in general and in learning, aligning with our instructional goals.
@@
Fifty-five students showed great interests in interaction with AI. More specifically, students appreciated the chances to ask questions, get responses and evaluate responses with newly learned prompting skills (e.g. ”I am interested in the different responses from the different question-asking approaches”). The learning design was also favored by students. Many of them gained great interest in the learning objectives: ”Knowing AI can be used to study or review coursework”, ”Great to know that AI can not only provide knowledge, but also create exercise questions.” Some students liked the in-time, comprehensive feedback design: ”(the platform) will give detailed instructions after checking the answers I reply.” The scenario design was also favored by 8 students (e.g. ”(I like) the scientific report of water cycle and want to know more about it”). Additionally, some students favored the visual elements of the learning platform, including the cartoon logos and its animations (e.g. ”I like the cute cartoon robot holding a sign to tell me whether my answer is correct”). In short, the parts that students favored highly overlap with our instructional goals.
@@
Students’ main challenges could be categorized into productive struggles (Warshauer 2015) which are essential for learning, and elements that increase their extraneous load that should be minimized. The most common productive struggle was experiencing difficulties in asking AI questions, which was the core skill we intended to teach in this activity. While students acknowledged their lack of competency in ”writing a prompt that leads to a more helpful answer”, many of their responses directly quoted some of the prompting guidelines in our instructional materials, which were strong indicators of students’ mastery of tips for writing better prompts. For instance, students admitted the difficulties in writing prompts by saying ”I didn’t write a concise question”, ”I didn’t provide context information”. These precise quotes from the instruction content in survey responses indicated students’ knowledge retention of prompting strategies, which served as evidence of our effective activity. Reasons for increasing extraneous load included limitations of AI and learning platform (e.g. ”It takes so long for AI to response”, ”I have trouble in logging in”), lack of variation on scenario content (e.g. ”All scenarios are in science subjects which I’m not interested in”), and noticeably, limitations of students’ basic computer skills (N = 22, e.g. ”I am not good at typing, it is tedious to type a lot to ask AI questions”). Some students struggled in typing, copying, and pasting information, which might hinder their prompt construction, even if they understood how to improve one.
Study 2: Assessment Iteration Follow-Up
Results from Study 1 revealed a ceiling effect in the pre-test, indicating that students in our study had a basic understanding to identify learning-oriented prompts. However, practice performance showed that they had not yet mastered writing such prompts effectively (Table 3). Therefore, to better measure their prompting literacy learning (Judson 2012; Staus, O’Connell, and Storksdieck 2021), we iterated the assessment and conducted Study 2 to evaluate the assessment quality. Study 2 utilized the same procedure and materials with a similar population, except for the iterated assessment questions. Study 2 focused on RQ4 about assessment questions:
@@
Iterated Assessment Design
@@
The lack of information about learners’ prior knowledge can reduce the effectiveness of MCQ distractors. In contrast, TF questions require students to make a decision on each item individually, offering more data points to assess both the learners’ true knowledge level and the quality of the question items, particularly at lower granularity (Brassil and Couch 2019). In addition, to overcome difficulties in forming high-quality distractors in MCQs, one approach is to collect learners’ common misconceptions as MCQ items. OE1, OE2, and OE3 are designed as a misconception collectors for LO-AI Capacity for Supporting Learning and LO2-Contexts to Use AI for Learning to facilitate further iterations on assessment questions.
The gap between high MCQ scores and poor prompt-writing performance from Study 1 suggested that LO-Effective Prompt Formation involves higher-order cognitive skills (falling under the analyze and create levels of Bloom’s Revised Taxonomy (Krathwohl 2002)), and developing MCQs that effectively assess these higher cognitive levels is challenging even for expert instructors in traditional disciplines (Douglas, Wilson, and Ennis 2012), let alone for an emerging topic. Therefore, we switched to open-ended questions and designed OE4 ((Given a prompt and the AI Chatbot response) Do you think the question can generate good learning material for your quiz preparation? Why or why not?) to assess students’ analyzing skill, and OE5 (Rewrite the question to AI to generate better preparation questions) to assess prompt-writing skill more directly.
@@
RQ4: Does the revised assessment better evaluate the targeted learning objectives?
We evaluated the assessment question quality using difficulty and discrimination indices, and reported the overall reliability using Cronbach’s Alpha. The comparison of the initial and revised assessment is used to answer RQ4. In particular, we compared the quality of these two assessments using the data collected from two studies and three widely-used assessment evaluation matrices: difficulty, discrimination, and reliability (Nitko 1996; Salkind 2017). Our analyses yielded two preliminary findings:
@@
The difficulty level and discrimination index are calculated based on the formula definitions provided in the previous work (Boopathiraj and Chellamani 2013). Figure 3 displays the difficulty level and discrimination index for all questions. A good discrimination index should be no less than 0.2 and ideally approach 1, while a good difficulty level falls within the range of [0.3, 0.7]. None of the MCQ questions fall into the desired range, while 60% of OE and 30% of TF questions satisfied the requirement, indicating an improvement on the item quality as assessment .
@@
With the initial success on difficulty level and discrimination index, Cronbach’s Alpha scores were also calculated to assess internal reliability of the assessment questions. The Cronbach’s Alpha analysis revealed moderate internal consistency for both versions of the assessment. The original version achieved a coefficient of 0.68, slightly higher than the iterated version’s 0.58. Although below the 0.70 benchmark for strong reliability, these values suggest that the items are reasonably cohesive given the small number of questions and participants. The lower reliability in the iterated version may reflect the inclusion of more diverse item types or targeted revisions aimed at specific constructs. Future large-scale administrations will be essential to verify whether these patterns persist and to refine items for stronger internal consistency.
Lessons Learned and Future Work
@@
First, our results provided design implications of teaching prompting literacy and AI. By experiencing the learning-by-doing prompting activities with LLM-generated immediate elaborated feedback in various scenarios, students gained more experience in including contexts in their prompts and increased confidence in prompt writing for learning. However, the lack of improvement in other aspects (e.g., conciseness, elaboration) and post-test performance may be attributed to dosage effects (Zhai et al. 2010), suggesting that more exposure is needed for aspects that students are less familiar with. For the task scenarios, survey responses indicated these increased interest and motivation. However, students expressed needs for more relatable and diverse subjects and scenarios to cater varied interests. We also acknowledged not comparing our approach with other AI literacy instruction methods. Future work should explore comparative studies or taking this study as a benchmark for further iterations.
@@
Lastly, this work contributed to the development of learning and assessment materials for prompting literacy, enriched by insights from our iterative design and data-driven refinement process. As the first local study on prompting literacy, learners’ baseline knowledge were unclear, thus iterative refinement through classroom-based deployments is essential to ensure the instructional material effectiveness. Expanding the variety and number of assessment questions can further help identify students’ knowledge gaps and misconceptions. Through these initial efforts, we established a basis for prompting literacy instruction. Future work should scale assessment materials to larger students with more items, enabling the application of models such as Item Response Theory (IRT) (Embretson and Reise 2000) to obtain statistically robust measures of assessment quality.

### v2

@@ 4.6 RQ3: How do students perceive the benefits and challenges of this experience?
Most learners (87%) reported that they learned AI-related knowledge, including how to use AI in learning (e.g. “I can’t ask AI to write answers directly, otherwise I won’t learn anything.”), how to ask AI questions effectively (e.g. “Adding background and context information in the prompt can help AI to form better answers”), and AI’s capabilities (e.g. “AI can help on exam preparation”, “AI also has problems that it doesnt́ know about”). These indicated learners’ gain on conceptual knowledge and skills of using AI in general and in learning, aligning with our instructional goals.
@@
Fifty-five students showed great interests in interaction with AI. More specifically, students appreciated the chances to ask questions, get responses and evaluate responses with newly learned prompting skills (e.g. “I am interested in the different responses from the different question-asking approaches”). The learning design was also favored by students. Many of them gained great interest in the learning objectives: “Knowing AI can be used to study or review coursework”, “Great to know that AI can not only provide knowledge, but also create exercise questions.” Some students liked the in-time, comprehensive feedback design: “(the platform) will give detailed instructions after checking the answers I reply.” The scenario design was also favored by 8 students (e.g. “(I like) the scientific report of water cycle and want to know more about it”). Additionally, some students favored the visual elements of the learning platform, including the cartoon logos and its animations (e.g. “I like the cute cartoon robot holding a sign to tell me whether my answer is correct”). In short, the parts that students favored highly overlap with our instructional goals.
@@
Students’ main challenges could be categorized into productive struggles (warshauer2015productive) which are essential for learning, and elements that increase their extraneous load that should be minimized. The most common productive struggle was experiencing difficulties in asking AI questions, which was the core skill we intended to teach in this activity. While students acknowledged their lack of competency in “writing a prompt that leads to a more helpful answer”, many of their responses directly quoted some of the prompting guidelines in our instructional materials, which were strong indicators of students’ mastery of tips for writing better prompts. For instance, students admitted the difficulties in writing prompts by saying “I didn’t write a concise question”, “I didn’t provide context information”. These precise quotes from the instruction content in survey responses indicated students’ knowledge retention of prompting strategies, which served as evidence of our effective activity. Reasons for increasing extraneous load included limitations of AI and learning platform (e.g. “It takes so long for AI to response”, “I have trouble in logging in”), lack of variation on scenario content (e.g. “All scenarios are in science subjects which I’m not interested in”), and noticeably, limitations of students’ basic computer skills (N = 22, e.g. “I am not good at typing, it is tedious to type a lot to ask AI questions”). Some students struggled in typing, copying, and pasting information, which might hinder their prompt construction, even if they understood how to improve one.
5 Study 2: Assessment Iteration Follow-Up
Results from Study 1 revealed a ceiling effect in the pre-test, indicating that students in our study had a basic understanding to identify learning-oriented prompts. However, practice performance showed that they had not yet mastered writing such prompts effectively (Table 3). Therefore, to better measure their prompting literacy learning (judson2012learning; staus2021addressing), we iterated the assessment and conducted Study 2 to evaluate the assessment quality. Study 2 utilized the same procedure and materials with a similar population, except for the iterated assessment questions. Study 2 focused on RQ4 about assessment questions:
@@ 5 Study 2: Assessment Iteration Follow-Up
5.1 Iterated Assessment Design
@@
The lack of information about learners’ prior knowledge can reduce the effectiveness of MCQ distractors. In contrast, TF questions require students to make a decision on each item individually, offering more data points to assess both the learners’ true knowledge level and the quality of the question items, particularly at lower granularity (brassil2019multiple). In addition, to overcome difficulties in forming high-quality distractors in MCQs, one approach is to collect learners’ common misconceptions as MCQ items. OE1, OE2, and OE3 are designed as a misconception collectors for LO-AI Capacity for Supporting Learning and LO2-Contexts to Use AI for Learning to facilitate further iterations on assessment questions.
The gap between high MCQ scores and poor prompt-writing performance from Study 1 suggested that LO-Effective Prompt Formation involves higher-order cognitive skills (falling under the analyze and create levels of Bloom’s Revised Taxonomy (krathwohl2002revision)), and developing MCQs that effectively assess these higher cognitive levels is challenging even for expert instructors in traditional disciplines (douglas2012multiple), let alone for an emerging topic. Therefore, we switched to open-ended questions and designed OE4 (Given a prompt and the AI Chatbot response, do you think the question can generate good learning material for your quiz preparation? Why or why not?) to assess students’ analyzing skill, and OE5 (Rewrite the question to AI to generate better preparation questions) to assess prompt-writing skill more directly.
@@
5.2 RQ4: Does the revised assessment better evaluate the targeted learning objectives?
We evaluated the assessment question quality using difficulty and discrimination indices, and reported the overall reliability using Cronbach’s Alpha. The comparison of the initial and revised assessment is used to answer RQ4. In particular, we compared the quality of these two assessments using the data collected from two studies and three widely-used assessment evaluation matrices: difficulty, discrimination, and reliability (nitko1996educational; salkind2017tests). Our analyses yielded two preliminary findings:
@@ 5.2 RQ4: Does the revised assessment better evaluate the targeted learning objectives?
The difficulty level and discrimination index are calculated based on the formula definitions provided in the previous work (boopathiraj2013analysis). Figure 3 displays the difficulty level and discrimination index for all questions. A good discrimination index should be no less than 0.2 and ideally approach 1, while a good difficulty level falls within the range of [0.3, 0.7]. None of the MCQ questions fall into the desired range, while 60% of OE and 30% of TF questions satisfied the requirement, indicating an improvement on the item quality as assessment .
@@
With the initial success on difficulty level and discrimination index, Cronbach’s Alpha analysis showed moderate internal reliability for both assessment versions: 0.68 for the original and 0.58 for the iterated version. Although slightly below the 0.70 benchmark, these values are reasonable given the small sample size and limited number of items. The lower reliability in the iterated version may reflect greater item diversity or more targeted revisions. Larger-scale administrations will be needed to confirm these patterns and further strengthen internal consistency.
6 Lessons Learned and Future Work
@@
First, our results provided design implications of teaching prompting literacy and AI. By experiencing the learning-by-doing prompting activities with LLM-generated immediate elaborated feedback in various scenarios, students gained more experience in including contexts in their prompts and increased confidence in prompt writing for learning. However, the lack of improvement in other aspects (e.g., conciseness, elaboration) and post-test performance may be attributed to dosage effects (zhai2010dosage), suggesting that more exposure is needed for aspects that students are less familiar with. Survey responses showed that task scenarios increased students’ interest and motivation, but students also wanted more relatable and diverse topics to match their varied interests. We also note that our study did not compare this approach to other AI literacy instructional methods; future work should conduct comparative studies or use this study as a benchmark for further iterations.
@@
Lastly, this work contributed to the development of learning and assessment materials for prompting literacy, informed by iterative design and data-driven refinement. As the first local study on prompting literacy, students’ baseline knowledge was unclear, making classroom-based iteration essential for ensuring instructional effectiveness. Increasing the variety and number of assessment items will help better surface learners’ misconceptions and knowledge gaps. These initial efforts establish a foundation for prompting literacy instruction. Future work should scale assessments to larger populations and items, using models such as Item Response Theory (IRT) (embretson2000irt) to produce statistically robust measures of assessment quality.